# Hyperfine van der Waals repulsion between open-shell polar molecules

### v1

@@ II Hyperfine van der Waals
From here on, we use CaF as our main example to illustrate hyperfine van der Waals interactions, although we expand on universality and other molecules in Sec. V. For CaF, the ground rotational state $j=0$ has two hyperfine states, namely $F=0$ and $1$, split by spin-spin coupling. The higher $j$ states have four levels due to additional spin-rotation coupling. For example, the hyperfine states for $j=1$ are given by $F=1^{-},0,1^{+},2$. Here, we label states of the same $F$ with -- and $+$ for those lower or higher in energy. In Figs. 2(a)-(c) the hyperfine structure is shown for rotational states $j=0$, 1 and 2 with energy levels split by around 25 to 123 MHz. In order to understand the second-order coupling, it is more insightful to look at the combined energy levels for the pair of molecules, which for $j+j^{\prime}=0+1$ and $1+2$ are given in Figs. 2(d) and 2(e), respectively. We discuss these separately below.
As highlighted in color in Fig. 2(d) for $j+j^{\prime}=0+1$, there exist two molecule-molecule states that do not interact through resonant dipole-dipole, but through hyperfine van der Waals. These are $20,F^{\prime}+1,F^{\prime}=0,0+1,$ 2 and $0,0+1,0$, shown in blue and red, respectively, since $0,0\to 1,2$ and $0,0\to 1,0$ are not dipole allowed. Whether the hyperfine van der Waals interaction in these states is attractive or repulsive depends on which other states they predominantly couple to. For $F+F^{\prime}=0+0$, this is the $F+F^{\prime}=1+1^{-}$ state, which is located higher in energy, resulting in an attractive van der Waals interaction. For $F+F^{\prime}=0+2$, the main coupling also occurs with $F+F^{\prime}=1+1^{-}$, which this time is located below in energy. This means that the hyperfine van der Waals interaction for $F+F^{\prime}=0+2$ is repulsive and that this state is the primary candidate within $j+j^{\prime}=0+1$ to find strongly suppressed loss rates.
Similar states can also be found within $j+j^{\prime}=1+2$ and so on, although the hyperfine structure is more involved, as seen in Fig. 2(e). Note that the number of hyperfine levels and the overall hyperfine structure remain the same for higher rotational states, as shown for $j+j^{\prime}=2+3$ in the Appendix. For $j+j^{\prime}=1+2$, the main candidate for a repulsive interaction is the $F+F^{\prime}=0+3$ state, shown in blue. The states that are shown in red either result in attractive van der Waals interactions, or are accidentally nearly degenerate with another state it couples to, still resulting in resonant dipole-dipole interactions. In general, for CaF, we find the most repulsive interactions between molecules in $j,F+(j+1),{F^{\prime}}$ when $F=j-1$ and $F^{\prime}=j+2$. This state can only couple to one other state, namely $F+F^{\prime}=j^{\prime}+j$, which is lower in energy.
@@ III Theoretical framework and computational details
with $C_{2}(\hat{R})$ the rank-2 tensor whose spherical components are the Racah normalized spherical harmonics, and
@@
×$\hskip 14.22636pt\times\bigg{[}\Big{[}C_{1}\left(\hat{r}^{(A)}\right)\otimes C_{1}\left(\hat{r}^{(B)}\right)\Big{]}^{(2)}\otimes C_{2}\left(\hat{R}\right)\bigg{]}^{(0)}_{0}\,,$ (7)
@@
×$\times\langle J,m+M_{S},I,M_{I}|F,M_{F}\rangle|j,m\rangle|S,M_{S}\rangle|I,M_{I}\rangle\,.$ (8)
@@
With the Hamiltonian and basis sets now discussed, we turn to collisions. The collisional loss rates are obtained through coupled-channels scattering calculations using the renormalized Numerov method [46]. The short-range losses are modeled by an absorbing boundary condition, which assumes 100% loss of flux going into the short-range reactive channels at the distance where this boundary condition is imposed. This matches experimentally observed losses for ultracold molecules [24, 25, 26, 27, 28, 29]. At large distances, we impose $S$-matrix boundary conditions [47]. With these boundary conditions, results are converged with a radial grid ranging from 50 to 100 000 $a_{0}$ with at least 30 points per local de Broglie wavelength. The resulting $S$-matrix after propagation contains the state-to-state scattering probabilities from which we get collisional cross sections, including those pertaining to collisional loss. There are two loss mechanisms: short-range loss through the absorbing boundary, and inelastic loss to other channels [39]. Inelastic collisions that change $M_{F}$ within the initial $F$ state are not considered as loss, as these states are degenerate at infinite separation and zero field. To obtain thermal rate coefficients, these energy-dependent cross sections are then averaged over a Maxwell-Boltzmann distribution, where we perform the integration using 37 logarithmically spaced collision energies between $10^{-2}$ and $10^{4}$ $\mu$K.
@@ IV Interaction potentials and loss rates
One other state also requires attention, the $F+F^{\prime}=0+1^{+}$ state. Although it has resonant first-order dipole-dipole interactions, it also couples to the $1+1^{-}$ state, which lies closely below by only 0.36 MHz $\approx$ 16 $\mu$K. This results in a repulsive potential that can decrease losses, but only when the collision energy is small compared to this 16 $\mu$K  [^1] The coincidental near degeneracy of $F+F^{\prime}=0+1^{+}$ and $1+1^{-}$ is sensitive to the precise hyperfine constants..
The distinct interactions among all hyperfine pairs lead to vastly different loss rates. For each hyperfine state of CaF within $j+j^{\prime}=0+1$, these rates are shown as a function of temperature in Fig. 4. The various hyperfine states with the highest loss rates correspond to resonant dipole-dipole losses. The $F+F^{\prime}=0+0$ state highlighted in red reaches slightly lower losses at low temperatures corresponding to universal loss due to its attractive $R^{-6}$ interaction [49]. The $0+1^{+}$ state shown in orange shows decreased loss rates at temperatures that are roughly lower than 16 $\mu$K. Finally, the state that has been of most interest so far, the $0+2$ state presented in blue, shows a loss rate that becomes as small as $10^{-14}$ cm3/s due to the potential barrier, which height is on the order of mK.
@@ V Universality and other diatomic molecules
Hyperfine van der Waals interactions are not unique to CaF. We find such interactions in even the simplest systems with a single spin $S=1/2$. Molecules can then be prepared in, e.g., $1,{1/2}+2,{5/2}$, where $\Delta F>1$ does not allow coupling to $2,{5/2}+1,{1/2}$. The dominant interaction then results from second-order dipole-dipole coupling to other $F+F^{\prime}$ states within $j+j^{\prime}=1+2$. Whether the resulting hyperfine van der Waals interaction are in general attractive or repulsive then depends on the specific molecular hyperfine structure.
[…]
@@
For the earth-alkaline monofluoride series, the hyperfine structure is rather similar, though the relative location of the energy levels does determine differences between them. At low temperatures, we find that loss rates increase as CaF $<$ SrF $<$ MgF $<$ BaF. For SrF and BaF, the main source of loss is short-range loss. As the coupling of the electron spin to the nuclear spin decreases (see $c_{4}$ in Table 1) and coupling to rotation increases ($c_{1}$ in Table 1), the gap between the $F+F^{\prime}=0+2$ and the $1+1^{-}$ state below increases, while the one to the $1+1^{+}$ state above decreases. This means that the repulsive barrier decreases from MgF to BaF. This increases short-range loss, which becomes dominant for SrF and BaF, hence why higher loss rates are expected than for CaF. For MgF and CaF, the residual loss is dominated by inelastic loss to other hyperfine states, which is relatively larger for MgF. The trade-off between short-range and inelastic losses results in the lowest loss rates for CaF.
@@ VI Dependence on magnetic fields
In zero field, each $F$ state has $12F+$ 1 degenerate $M_{F}$ components. So far, we have assumed $M_{F}$-changing collisions within $F$ do not constitute loss. Magnetic fields, however, split up these $M_{F}$ states by the Zeeman effect. Collisions that change $M_{F}$ could then lead to energy release and therefore loss or heating, depending on the trap depth. We therefore explore for CaF up to what magnetic field strength collisional loss is still suppressed, where we now consider $M_{F}$-changing collisions to constitute loss due to the energy released by Zeeman relaxation.
@@
with electronic $g$-factor $g_{e}\approx 2.002319$ and bohr magneton $\mu_{B}=1/2$ in atomic units. The coupling of the magnetic field with the nuclear spin or the rotation of the molecule is neglected. We add the Zeeman Hamiltonian of Eq. (9) to the monomer Hamiltonian (3).
For the $F+F^{\prime}=0+2$ state, the molecule in $F=2$ has components $M_{F}=-2,-1,0,1,2$, with $M_{F}=-2$ the lowest in energy in a magnetic field. For CaF, the energy levels at infinite separation as a function of magnetic field are given in Fig. 8(a), where it can be seen that states start to cross above magnetic fields of around 10 G. The loss rates of these states at $T=1$ $\mu$K are shown in Fig. 8(b). At small $B$, the inelastic loss to other $F+F^{\prime}$ states is around $10^{-13}$ cm3/s, as shown by the dashed curve. At large $B$, where $F+F^{\prime}$ levels cross, the loss rate increases and is dominated by transitions to other $F+F^{\prime}$ states. For fields below 10 G, it is the lowest $M_{F}=-2$ state that shows the lowest loss rates. For molecules in other $M_{F}$ states, the dominant loss is Zeeman relaxation to other $M_{F}$ states within $F+F^{\prime}=0+2$, with energy released due to the Zeeman splitting. Below 0.1 G, this splitting becomes smaller than the collision energy, where excitation to other $M_{F}$ becomes energetically available. This also means that the energy release has become small enough that it might not necessarily lead to heating. In general, we find that the suppression of loss rates due to hyperfine van der Waals repulsion should be observable for finite magnetic fields below 10 G.
Figure 8: Magnetic field dependence. (a) Energy levels of both molecules at infinite separation as a function of magnetic field $B$. (b) Loss rates as a function of magnetic field at $T=1$ $\mu$K. The different curves correspond to the $F=2$ molecule prepared in its different $M_{F}$ components. For $M_{F}=-2$, the additional dashed curve shows loss rates that exclude excitation to other $M_{F}$ states at low $B$.
@@ VII Conclusion
The CaF hyperfine energy level structure is very similar for any $j\geq 1$, see Fig. 2; the number of hyperfine states and their positions are the same, while the splitting due to spin-rotation coupling gradually increases with $j$. As a result, the hyperfine energy levels for a pair of CaF molecules in $j+(j+1)$ with $j\geq 1$ are also very similar for different $j$. In Fig. 9, we give the CaF hyperfine structure for the combined pair of $j+j^{\prime}=2+3$. The similarities are clear when comparing to $j+j^{\prime}=1+2$ in Fig. 2(e), which shows that the relative positions can change slightly, while the overall structure remains. The same principle holds for higher $j$. In general, for CaF, we find the most repulsive interactions between molecules in $j,F+(j+1),{F^{\prime}}$ when $F=j-1$ and $F^{\prime}=j+2$.

### v2

@@ II Hyperfine van der Waals
From here on, we use CaF as our main example to illustrate hyperfine van der Waals interactions, although we expand on universality and other molecules in Sec. V. For CaF, the ground rotational state $j=0$ has two hyperfine states, namely $F=0$ and $1$, split by spin-spin coupling. The higher $j$ states have four levels due to additional spin-rotation coupling. For example, the hyperfine states for $j=1$ are given by $F=1^{-},0,1^{+},2$. Here, we label states of the same $F$ with - and ++ for those lower or higher in energy. In Figs. 2(a)-(c) the hyperfine structure is shown for rotational states $j=0$, 1 and 2 with energy levels split by around 25 to 123 MHz. The hyperfine splittings are determined by diagonalization of the hyperfine Hamiltonian in Eq. 3 with coupling parameters as given in Table 1. In order to understand the second-order coupling, it is more insightful to look at the combined energy levels for the pair of molecules, which for $j+j^{\prime}=0+1$ and $1+2$ are given in Figs. 2(d) and 2(e), respectively. We discuss these separately below.
As highlighted in color in Fig. 2(d) for $j+j^{\prime}=0+1$, there exist two molecule-molecule states that do not interact through resonant dipole-dipole, but through hyperfine van der Waals. These are $0,F^{\prime}+1,F^{\prime}=0,0+1,2$ and $0,0+1,0$, shown in blue and red, respectively, since $0,0\to 1,2$ and $0,0\to 1,0$ are not dipole allowed. Whether the hyperfine van der Waals interaction in these states is attractive or repulsive depends on which other states they predominantly couple to. For $F+F^{\prime}=0+0$, this is the $F+F^{\prime}=1+1^{-}$ state, which is located higher in energy, resulting in an attractive van der Waals interaction. For $F+F^{\prime}=0+2$, the main coupling also occurs with $F+F^{\prime}=1+1^{-}$, which this time is located below in energy. This means that the hyperfine van der Waals interaction for $F+F^{\prime}=0+2$ is repulsive and that this state is the primary candidate within $j+j^{\prime}=0+1$ to find strongly suppressed loss rates.
Similar states can also be found within $j+j^{\prime}=1+2$ and so on, although the hyperfine structure is more involved, as seen in Fig. 2(e). Note that the number of hyperfine levels and the overall hyperfine structure remain the same for higher rotational states, as shown for $j+j^{\prime}=2+3$ in the Appendix. For $j+j^{\prime}=1+2$, the main candidate for a repulsive interaction is the $F+F^{\prime}=0+3$ state, shown in blue. The states that are shown in red either result in attractive van der Waals interactions, or are accidentally nearly degenerate with another state it couples to, still resulting in resonant dipole-dipole interactions. In general, for CaF, we find the most repulsive interactions between molecules in $j,F+(j+1),{F^{\prime}}$ when F=j1F=j-1 and $F^{\prime}=j+2$. This state can only couple to one other state, namely $F+F^{\prime}=j^{\prime}+j$, which is lower in energy.
@@ III Theoretical framework and computational details
[…]
@@
$\hskip 14.22636pt\times\bigg[\Big[C_{1}\left(\hat{r}^{(A)}\right)\otimes C_{1}\left(\hat{r}^{(B)}\right)\Big]^{(2)}\otimes C_{2}\left(\hat{R}\right)\bigg]^{(0)}_{0}\,$, (7)
@@
$\times\langle J,m+M_{S},I,M_{I}|F,M_{F}\rangle|j,m\rangle|S,M_{S}\rangle|I,M_{I}\rangle\,.$ (8)
@@
[…]
@@ IV Interaction potentials and loss rates
[…]
@@ V Universality and other diatomic molecules
Hyperfine van der Waals interactions are not unique to CaF. We find such interactions in even the simplest systems with a single spin $S=1/2$. Molecules can then be prepared in, e.g., 1,1/2+2,5/21,{1/2}+2,{5/2}, where $\Delta F>1$ does not allow coupling to 2,5/2+1,1/22,{5/2}+1,{1/2}. The dominant interaction then results from second-order dipole-dipole coupling to other $F+F^{\prime}$ states within $j+j^{\prime}=1+2$. Whether the resulting hyperfine van der Waals interaction are in general attractive or repulsive then depends on the specific molecular hyperfine structure.
Hyperfine van der Waals repulsion depends on the magnitude of the hyperfine energy scale compared to the dipolar energy scale. We show this in Fig. 6, where we artificially scale all the hyperfine terms in the CaF Hamiltonian (3) by an overall scaling parameter $\lambda$ such that relative splittings are kept identical. At larger overall hyperfine coupling, the separation between energy levels increases. Since the repulsive interaction potential forms a barrier due to coupling with higher-lying levels, the barrier height increases as well. This effect therefore decreases loss rates at larger hyperfine coupling. Scaling the hyperfine couplings up by more than a factor of $\lambda=10^{2}$ causes the hyperfine energy levels of $j+j^{\prime}=0+1$ to cross those of the $j+j^{\prime}=1+2$ manifold. This shows that for hyperfine couplings smaller than the rotational constant, which is typically the case, larger hyperfine splittings lead to lower loss rates. Scaling down the hyperfine couplings leads to loss rates similar to resonant dipole-dipole collisions, as if there were no hyperfine.
@@
For the earth-alkaline monofluoride series, the hyperfine structure is rather similar, though the relative location of the energy levels does determine differences between them. At low temperatures, we find that loss rates increase as CaF << SrF << MgF << BaF. For SrF and BaF, the main source of loss is short-range loss. As the coupling of the electron spin to the nuclear spin decreases (see $c_{4}$ in Table 1) and coupling to rotation increases ($c_{1}$ in Table 1), the gap between the $F+F^{\prime}=0+2$ and the $1+1^{-}$ state below increases, while the one to the $1+1^{+}$ state above decreases. This means that the repulsive barrier decreases from MgF to BaF. This increases short-range loss, which becomes dominant for SrF and BaF, hence why higher loss rates are expected than for CaF. For MgF and CaF, the residual loss is dominated by inelastic loss to other hyperfine states, which is relatively larger for MgF. The trade-off between short-range and inelastic losses results in the lowest loss rates for CaF.
@@ VI Dependence on magnetic fields
In zero field, each $F$ state has $2F+1$ degenerate $M_{F}$ components. So far, we have assumed $M_{F}$-changing collisions within $F$ do not constitute loss. Magnetic fields, however, split up these $M_{F}$ states by the Zeeman effect. Collisions that change $M_{F}$ could then lead to energy release and therefore loss or heating, depending on the trap depth. We therefore explore for CaF up to what magnetic field strength collisional loss is still suppressed, where we now consider $M_{F}$-changing collisions to constitute loss due to the energy released by Zeeman relaxation.
@@
with electronic $g$-factor $g_{e}\approx 2.002319$ and bohr magneton $\mu_{B}=1/2$ in atomic units. The coupling of the magnetic field with the nuclear spin or the rotation of the molecule is neglected, as well as the anisotropic coupling with the electron spin. These further contributions are at least three orders of magnitude smaller than the isotropic electron spin interaction we include here [61]. We add the Zeeman Hamiltonian of Eq. (9) to the monomer Hamiltonian (3).
[…]
@@ VII Conclusion
The CaF hyperfine energy level structure is very similar for any $j\geq 1$, see Fig. 2; the number of hyperfine states and their positions are the same, while the splitting due to spin-rotation coupling gradually increases with $j$. As a result, the hyperfine energy levels for a pair of CaF molecules in $j+(j+1)$ with $j\geq 1$ are also very similar for different $j$. In Fig. 9, we give the CaF hyperfine structure for the combined pair of $j+j^{\prime}=2+3$. The similarities are clear when comparing to $j+j^{\prime}=1+2$ in Fig. 2(e), which shows that the relative positions can change slightly, while the overall structure remains. The same principle holds for higher $j$. In general, for CaF, we find the most repulsive interactions between molecules in $j,F+(j+1),{F^{\prime}}$ when F=j1F=j-1 and $F^{\prime}=j+2$.